The following financial information was obtained from the year ended 2018 income statements for Cash Automotive and Penningto$ Net income Income tax expense Interest expense Cash Pennington 26,070 $ 74,188 9,270 27,080 300 2,900
Answer #1

First Understand the Time interest Earned ration Concepts and then solving this problem

Times interest earned (TIE) or interest coverage ratio is a measure of a company's ability to honor its debt payments. It may be calculated as either EBIT or EBITDA divided by the total interest expense.

{\displaystyle {\mbox{Times-Interest-Earned}}={\frac {\mbox{EBIT or EBITDA}}{\mbox{Interest Expense}}}}

When the interest coverage ratio is smaller than one, the company is not generating enough cash from its operations EBIT to meet its interest obligations. The Company would then have to either use cash on hand to make up the difference or borrow funds. Typically, it is a warning sign when interest coverage falls below 2.5x.

The time's interest earned ratio indicates the extent to which earnings are available to meet interest payments.

A lower times interest earned ratio means less earnings are available to meet interest payments and that the business is more vulnerable to increases in interest rates and being unable to meet their existing outstanding loan obligations.

As Per Data given above case first Company ratio is

Times-interest Earned = EBIT/Interest expenses

EBIT = Net income minus Interest and Taxes

= 26070-9270-300 =16500

TIER = 16500/300 = 55

for Second Company

= 74188-27080-2900=44208

TIER = 44208/2900= 15.24

Ans = 15.24 for Pennington Automotive

So as Compere above two ratio Cash Automotive company Better able recovered their interest expenses